OpenStudy (anonymous):

The measure of an angle is three more than twice its supplement. What is the measure of the angle?

OpenStudy (anonymous):

@Hero

hero (hero):

The sum of an angle and its supplement is 180 degrees. \(\angle{1} + \angle{2} = 180\)

hero (hero):

Let: \(\angle{1} = x\) \(\angle{2} = 2x + 3\)

OpenStudy (anonymous):

Let x = smaller angle Let 2x+6 be the larger angle The angles are complementary, so their sum is 90° Therefore.... x + (2x+6) = 90 3x + 6 = 90 3x = 84 x = 84/3 = 28 So the smaller angle measures 28° and the larger angle measures 62°; as a check, note that they add up to 90° ANSWER 28° and 62°

hero (hero):

Where'd the 6 in 2x + 6 come from?

hero (hero):

@trina88

OpenStudy (anonymous):

is it supposed to be 3 instead of 6?

hero (hero):

Of course. For example Suppose you are 15 years old and suppose your mom's age is 3 more than twice your age, then she'd be 33 years old.

OpenStudy (anonymous):

i see, it makes sense... so was @trina88 right with the equation part? can i just plug in 3 instead of 6?

hero (hero):

Supplementary and complementary do not mean the same thing.